OpenStudy (anonymous):

I am trying to figure out an appropriate method in determining how many terms there are for any given itteration So basically We are given an example where any particular stock can either go up by a factor or down by a factor after every closing period. So for instance after 5 closing periods how many term will there be taking into consideration that order obv doesn't matter since UUD=DUU=UDU

OpenStudy (anonymous):

I believe the number you're looking for is perm(x)/x! Does that sound familiar?

OpenStudy (anonymous):

what does perm(x) mean?

OpenStudy (anonymous):

ohhh permutation

OpenStudy (anonymous):

Or wait, \[\frac{\left(\begin{matrix}n \\ 2\end{matrix}\right)}{n!}\]

OpenStudy (anonymous):

Right? I think dividing by n! covers the issue with order.

OpenStudy (anonymous):

Haha whatever I'm doing is way off, but conceptually close I think. Obviously my number is less than 1.

OpenStudy (anonymous):

ya no that wont work lemme draw a graph for the 3rd iteration

OpenStudy (anonymous):



OpenStudy (anonymous):

S represents the stock u=up factor d=down factor

OpenStudy (anonymous):

So for the third iteration you have 4 choices

OpenStudy (anonymous):

Could it be \[\frac{2^n}{(n-1)!}\]

OpenStudy (anonymous):

hmmmmmmm

OpenStudy (anonymous):

No, that's still not right for the 2nd iteration. It's right for 1 and 3 though. Weird.

OpenStudy (anonymous):

kai you lil bish ...

OpenStudy (anonymous):

So it's definitely going to be 2^n minus something. What you have is n spots, each with two options. Hence 2^n. But because of repetition, you need a way to count those that are repeated.

OpenStudy (kainui):

lol I'm not totally sure what you're asking for but like there will always be only n+1 possible choices and they will be distributed like @escolas is saying.

OpenStudy (anonymous):

Haha wow, right..

OpenStudy (anonymous):

omg -.-

OpenStudy (anonymous):

this is embarrassing

OpenStudy (kainui):

Hahahaha this is called "Pascal's Triangle" lol

OpenStudy (anonymous):

hahaha

OpenStudy (anonymous):

This isnt even funny ,,, this is just sad

OpenStudy (anonymous):

Can I erase all my embarrassing answers?

OpenStudy (anonymous):

thanks escolas for attempting to help :)

OpenStudy (anonymous):

ya sure :)

OpenStudy (anonymous):

Hahah any time.. :P

OpenStudy (anonymous):

Haha not sure I deserved that medal, but thank you very much. Cheers.